# TwinVLA: Data-Efficient Bimanual Manipulation with Twin Single-Arm Vision-Language-Action Models

### v1

@@ A.3 SingleVLA VLM Ablation
We validate SingleVLA’s VLM choice in the LIBERO (liu2023libero) environment using several VLMs. The LIBERO actions are converted to absolute EEF 6D control. Due to computational limits, we directly fine-tune the pretrained VLM checkpoints on LIBERO (i.e., without additional pretraining on LIBERO). Each model is evaluated with $500$ rollouts per task suite under identical random seeds. Results are shown in Table 3.
@@
Qwen2VL-2B (wang2024qwen2vlenhancingvisionlanguagemodels) $0.4\% 88$$8.6\%$ $83.8\%$ $43.0\%$ $73.9\%$
InternVL2.5-1B (chen2025expandingperformanceboundariesopensource) $64.6\%$ $84.8\%$ $78.4\%$ $46.2\%$ $68.5\%$
Eagle2-1B (li2025eagle2buildingposttraining) $73.4\%$ $85.4\%$ $90.8\%$ $46.6\%$ $\mathbf{74.0\%}$
@@
OpenVLA (kim2024openvla) $4.7\% 88$$8.4\%$ $79.2\%$ $53.7\%$ $76.5\%$
Octo (octo_2023) $78.9\%$ $85.7\%$ $84.6\%$ $51.1\%$ $75.1\%$
@@
Appendix B Implementation Details
@@
We use the official implementation of RDT-1B. Diffusion Policy and $\pi_{0}$ are evaluated via the public LeRobot release (cadene2024lerobot), with two modifications for a fair comparison. First, the LeRobot evaluation script normalized images differently from training; we corrected this to match the training pipeline.
@@ C.1 Joint Attention
The joint attention in TwinVLA is fundamentally almost identical to the implementation in the Mixture-of-Transformers (MoT) (liang2024mixtureoftransformerssparsescalablearchitecture) study. While MoT has transformers for text, image, and speech inputs, in TwinVLA, the inputs for the left and right arms correspond to these.
Furthermore, MoT requires an operation to group mixed inputs by modality and then restore their original order. However, this process is unnecessary in TwinVLA because the inputs are fed in a fixed sequence: left arm, then right arm. The detailed computation process is shown in Algorithm˜1.
Algorithm 1 Joint Attention Computation
1:Let $x=(x_{1},\dots,x_{n})$ be the input sequence with $x_{i}\in\mathbb{R}^{d}$; let $m_{i}\in\{\text{left arm},\text{right arm}\}$ denote the modality of $x_{i}$.
2:Let $\mathcal{M}=\{\text{left arm},\text{right arm}\}$.
3:for each modality $m\in\mathcal{M}$ do
4:  $I_{m}\leftarrow\{i:m_{i}=m\}$ $\triangleright$ Indices for modality $m$
5:  $X_{m}\leftarrow\{x_{i}:i\in I_{m}\}$ $\triangleright$ Group tokens by modality
6:  $Q_{m}\leftarrow W_{Q}^{m}X_{m},\;K_{m}\leftarrow W_{K}^{m}X_{m},\;V_{m}\leftarrow W_{V}^{m}X_{m}$ $\triangleright$ Modality-specific projections
7:end for
[…]
10:for each modality $m\in\mathcal{M}$ do
11:  $O_{m}\leftarrow W_{O}^{m}A|_{I_{m}}$ $\triangleright$ Modality-specific output projection
12:  $H_{m}\leftarrow X_{m}+\mathrm{LayerNorm}^{m}_{\text{attn}}(O_{m})$ $\triangleright$ Residual & norm
13:  $F_{m}\leftarrow\mathrm{FFN}_{m}(H_{m})$
14:  $Y_{m}\leftarrow H_{m}+\mathrm{LayerNorm}^{m}_{\text{ffn}}(F_{m})$
15:end for
16:return $\{Y_{m}:m\in\mathcal{M}\}$
C.2 Attention Re-weighting
@@
This discrepancy requires greater adaptation effort for TwinVLA during fine-tuning on bimanual tasks. To address this, we introduce a simple re-weighting trick immediately after the attention scores are calculated. Specifically, we double the attention weights corresponding to the shared modality tokens and then re-normalize all weights to sum to one. This adjustment effectively restores the proportional attention each VLM backbone assigns to the shared inputs, aligning it with the pretraining conditions (see Figure˜11). Applying this method reduced the initial fine-tuning loss by approximately 40%. While TwinVLA could learn bimanual manipulation without this technique, the required adaptation period would be substantially longer. This simple trick makes the process significantly more efficient and faster. We illustrate our implementation with simple pseudocode in Algorithm˜2.
Algorithm 2 Attention Re-weighting
Input: Attention weights $\mathbf{A}$, Scale factor $\alpha=2$
2:Output: Re-weighted attention weights $\mathbf{A}^{\prime}$
function ApplyReweighting($\mathbf{A}$, $\mathbf{c}$, $\alpha$)
4:  Create mask $\mathbf{M}$ $\triangleright$ Create a mask for arm-specific inputs
  $\mathbf{A_{reweighted}}\leftarrow\mathbf{A}\odot(\mathbf{M}+\alpha\cdot\neg\mathbf{M})$ $\triangleright$ Apply scaling to attention weights using the mask
6:  $\mathbf{A_{reweighted}}\leftarrow\text{Normalize}(\mathbf{A_{reweighted}})$ $\triangleright$ Normalize the new weights
8:end function
C.3 MoE integration
Feed-Forward Networks. To share FFNs, we adopt the common approach of using a gating-based MoE. In standard MoE, multiple FFNs are included within a transformer, and a gating mechanism activates a subset for each input. In TwinVLA, the two VLMs act as distinct FFN experts. Because shared inputs (e.g., egocentric views or language prompts) may have asymmetric relevance for each arm, the gating mechanism learns how much each FFN should contribute to processing the shared input. This approach is widely used and has been shown to improve training stability and preserve information more effectively than simple averaging (shazeer2017outrageously).
Other Components. Beyond FFNs, elements such as layer normalization and projection layers also require integration. For these, we apply task arithmetic (10.5555/3692070.3694018), merging the two VLMs via simple parameter averaging with weight $\lambda=0.5$. This extends MoE-style computation to the full transformer architecture.
@@ D.2 Quantitative Results
Table 6: Success rates for each model across all subtasks. The best overall performance is highlighted in bold. As $\pi_{0}$ is included as a skyline, as this is excluded from this direct comparison.
@@
Entirely off $0.20$ $0.55$ $0.60$ 0.65
We provide the quantitative results on real-world experiments in subtask-level detail in Table˜6. The results reveal the main bottleneck in each long-horizon task. The Carrot to bag task is challenging when inserting the carrot, which requires precisely opening the bag. The Brush to dustpan task’s bottleneck is the high-precision insertion of the brush into the dustpan. Lastly, in Take towel off rack, the final unfolding is difficult—unlike the simple initial steps—as it requires a successful switch between the arms. In the next subsection, we show qualitative results from these specific bottleneck phases.
@@ D.3 Qualitative Results
Figure˜13 presents qualitative results highlighting challenging situations for each task. A check mark was used when the model succeeded with a probability above 0.5, an X mark for probabilities below 0.3, and an exclamation mark icon for intermediate cases.
@@ D.4 Robot Hardware Spec
We conduct our real-world experiments using a custom-built robot named Anubis. The platform features a teleoperation system inspired by the Mobile ALOHA setup (fu2024mobile). Each arm has 6 DoF and is equipped with a parallel gripper and a wrist-mounted camera. At the center of the robot, an Intel RealSense camera is mounted on a height-adjustable mechanism, serving as the ego-centric view camera. Details are described in Table˜7. Anubis is equipped with a 3-wheel omni-directional base that supports planar locomotion; however, in this work, the mobility feature is not utilized.
@@ E.1 Tabletop-Sim
To test bimanual policies in simulation, we developed Tabletop-Sim, a new benchmark specifically engineered to evaluate dexterous manipulation skills, in contrast to other benchmarks (mu2025robotwindualarmrobotbenchmark) that primarily focus on task diversity. The benchmark comprises four single-task environments and one multi-task setup. Our task selection was guided by the taxonomy in DexMimicGen (11127809), which categorizes bimanual tasks into: (1) parallel (two arms are doing separate tasks simultaneously), (2) coordinated (two arms are closely working together), and (3) sequential (one arm completes the task, and the other arm takes over) interactions. Using a custom controller similar to GELLO (wu2024gellogenerallowcostintuitive), we collected 50 demonstrations for each single-task and 60 for the multi-task environment.
@@
Furthermore, to evaluate policy robustness, we established two difficulty settings for the four single-tasks. The original tasks are designated as the Easy setting, while a Hard variant for each task incorporates challenging variations such as different textures, object models, and the presence of distractor objects. Figure˜15 presents snapshots of each task.
To ensure reproducibility and support future research, we will fully open-source this simulation and dataset.
@@ E.2 Quantitative Results
This section describes the detailed results for the simulation tasks. The results for Tabletop-Sim are listed in  Table˜8, while the results for the RoboTwin 2.0 benchmark are in  Table˜9. For RoboTwin, the results for other baselines were referenced from the official benchmark results.

### v2

@@ A.3 SingleVLA VLM Ablation
We validate SingleVLA’s VLM choice in the LIBERO (Liu et al., 2023a) environment using several VLMs. The LIBERO actions are converted to absolute EEF 6D control. Due to computational limits, we directly fine-tune the pretrained VLM checkpoints on LIBERO (i.e., without additional pretraining on LIBERO). Each model is evaluated with $500$ rollouts per task suite under identical random seeds. Results are shown in Table 4.
@@
Qwen2VL-2B (Wang et al., 2024) $0.4\% 88$$8.6\%$ $83.8\%$ $43.0\%$ $73.9\%$
InternVL2.5-1B (Chen et al., 2025b) $64.6\%$ $84.8\%$ $78.4\%$ $46.2\%$ $68.5\%$
Eagle2-1B (Li et al., 2025) $73.4\%$ $85.4\%$ $90.8\%$ $46.6\%$ $\mathbf{74.0\%}$
@@
OpenVLA (Kim et al., 2024) $4.7\% 88$$8.4\%$ $79.2\%$ $53.7\%$ $76.5\%$
Octo (Octo Model Team et al., 2024) $78.9\%$ $85.7\%$ $84.6\%$ $51.1\%$ $75.1\%$
@@
Appendix B Training Details
@@
We use the official implementation of RDT-1B. Diffusion Policy and $\pi_{0}$ are evaluated via the public LeRobot release (Cadene et al., 2024), with two modifications for a fair comparison. First, the LeRobot evaluation script normalized images differently from training; we corrected this to match the training pipeline.
@@ Appendix C TwinVLA Details
Algorithm 2 Joint Attention
1:function JointAttention($\{Q_{m}\},\{K_{m}\},\{V_{m}\},M$)
2:  $Q,K,V\leftarrow\text{Concatenate}(\{Q_{m}\},\{K_{m}\},\{V_{m}\})$ $\triangleright$ Concatenate modality-specific Q, K, V
3:  $S\leftarrow\mathrm{Softmax}((QK^{\top}/\sqrt{d_{k}})+M)$ $\triangleright$ Apply causal joint mask M (Figure˜3(a))
4:  $S\leftarrow\text{\color[rgb]{0.078125,0.4375,0.79296875}ApplyReweighting}(S)$ $\triangleright$ Apply re-weighting (Algorithm˜4)
5:  $A\leftarrow S\cdot V$ $\triangleright$ Calculate output A
6:  return $\{A_{m}\}\leftarrow\text{Split}(A)$ $\triangleright$ Split output AA into modality-specific $A_{m}$
7:end function
@@ C.1 Joint Attention
The joint attention in TwinVLA is fundamentally almost identical to the implementation in the Mixture-of-Transformers (MoT) (Liang et al., 2024), but we applied attention-reweighting (Section˜C.3). While MoT has transformers for text, image, and speech inputs, in TwinVLA, the inputs for the left and right arms correspond to these.
Furthermore, MoT requires an operation to group mixed inputs by modality and then restore their original order. However, this process is unnecessary in TwinVLA because the inputs are fed in a fixed sequence: left arm, then right arm. The detailed computation process is shown in Algorithm˜2.
C.2 MoE integration
Feed-Forward Networks. To share FFNs, we adopt the common approach of using a gating-based MoE. In standard MoE, multiple FFNs are included within a transformer, and a gating mechanism activates a subset for each input. In TwinVLA, the two VLMs act as distinct FFN experts.
Because shared inputs (e.g., ego-centric views or language prompts) may have asymmetric relevance for each arm, the gating mechanism learns how much each FFN should contribute to processing the shared input. This approach is widely used and has been shown to improve training stability and preserve information more effectively than simple averaging (Shazeer et al., 2017). We computed $w_{\text{left}}$ by applying a simple linear layer and softmax to the token embeddings.
Other Components. Beyond FFNs, elements such as layer normalization and projection layers also require integration. For these, we apply task arithmetic (Tang et al., 2024), merging the two VLMs via simple parameter averaging with weight $\lambda=0.5$, elaborated Algorithm˜3. This extends MoE-style computation to the full transformer architecture.
Algorithm 3 Integration of other components
1:Let $\text{Projection}_{b}$ be projection layer from each backbone $b\in\{\text{left},\text{right}\}$.
2:Let $\text{LayerNorm}_{b}$ be layernorm from each backbone $b\in\{\text{left},\text{right}\}$.
3:function Proj($X^{m}$)
4:  if $m=\text{shared}$ then
5:   $F^{m}\leftarrow 0.5\cdot(\text{Projection}_{\text{left}}(X^{m})+\text{Projection}_{\text{right}}(X^{m}))$ $\triangleright$ Task arithmetic
6:  else
7:   $F^{m}\leftarrow\text{Projection}_{m}(X^{m})$
8:  end if
9:  return $F^{m}$
10:end function
11:
12:function Norm($X^{m}$)
13:  if $m=\text{shared}$ then
14:   $F^{m}\leftarrow 0.5\cdot(\text{LayerNorm}_{\text{left}}(X^{m})+\text{LayerNorm}_{\text{right}}(X^{m}))$ $\triangleright$ Task arithmetic
15:  else
16:   $F^{m}\leftarrow\text{LayerNorm}_{m}(X^{m})$
17:  end if
18:  return $F^{m}$
19:end function
C.3 Attention Re-weighting
@@
Algorithm 4 Attention Re-weighting
function ApplyReweighting($\mathbf{A}$, $\alpha=2$)
2:  Create mask $\mathbf{M_{r}}=(m\not=\text{shared})$ $\triangleright$ Create a mask for arm-specific inputs
  $\mathbf{A_{reweighted}}\leftarrow\mathbf{A}\odot(\mathbf{M_{r}}+\alpha\cdot\neg\mathbf{M_{r}})$ $\triangleright$ Apply scaling to attention weights using the mask
4:  $\mathbf{A_{reweighted}}\leftarrow\text{Normalize}(\mathbf{A_{reweighted}})$ $\triangleright$ Normalize the new weights
6:end function
@@
This discrepancy requires greater adaptation effort for TwinVLA during fine-tuning on bimanual tasks. To address this, we introduce a simple re-weighting trick immediately after the attention scores are calculated. Specifically, we double the attention weights corresponding to the shared modality tokens and then re-normalize all weights to sum to one. This adjustment effectively restores the proportional attention each VLM backbone assigns to the shared inputs, aligning it with the pretraining conditions (see Figure˜11). Applying this method reduced the initial fine-tuning loss by approximately 40%. While TwinVLA could learn bimanual manipulation without this technique, the required adaptation period would be substantially longer. This simple trick makes the process significantly more efficient and faster. We illustrate our implementation with simple pseudocode in Algorithm˜4.
@@ D.1 Task details
Figure 13: Challenging scene of Fold towel task.
Furthermore, to evaluate policy robustness in the real world, we tested the Fold towel task under more challenging conditions, such as reduced lighting and the presence of distractors. These scenarios are visualized in Figure˜13.
@@ D.2 Quantitative Results
Table 7: Success rates for each model across all subtasks. The best overall performance is highlighted in bold. As $\pi_{0}$ is included as an upper-bound, as this is excluded from this direct comparison.
@@
Fold towel First fold $0.00$ $1.00$ $0.90$ $1.00$
Rotate $0.00$ $1.00$ $0.85$ $1.00$
Second fold $0.00$ 0.90 $0.45$ 0.90
Extract hexkey Pick up $0.60$ $0.90$ $0.90$ $1.00$
Extract $0.35$ $0.80$ $0.55$ $0.90$
Put into bowl $0.30$ 0.80 $0.45$ 0.80
@@
Entirely off $0.20$ $0.45$ $0.60$ 0.65
We provide the quantitative results on real-world experiments in subtask-level detail in Table˜7. The results reveal the main bottleneck in each long-horizon task. First, for the two tasks, Fold towel and Extract hexkey, requiring tightly coupled bimanual coordination, the phase where both arms meet to execute the action appears to be critical. The Carrot to bag task is challenging when inserting the carrot, which requires precisely opening the bag. The Brush to dustpan task’s bottleneck is the high-precision insertion of the brush into the dustpan. Lastly, in Take towel off, the final unfolding is difficult—unlike the simple initial steps—as it requires a successful switch between the arms. In the next subsection, we show qualitative results from these specific bottleneck phases.
@@ D.3 Qualitative Results
Figure˜14 presents qualitative results highlighting challenging situations for each task. A check mark was used when the model succeeded with a probability above 0.5, an X mark for probabilities below 0.3, and an exclamation mark icon for intermediate cases.
@@
  • Fold towel. $\pi_{0}$ and TwinVLA successfully completed the task. RDT also generally performed well, but occasionally failed to fully rotate the towel by 90 degrees, which caused downstream failures. DP experienced substantial difficulty with the fold-towel task and ultimately failed to solve it.
  • Extract hexkey. $\pi_{0}$ and TwinVLA generally solved the task reliably. RDT performed the subtask of lifting the hexkey case well but often failed during extraction due to insufficient precision in grasping the hexkey once the case was lifted. DP failed both to reliably pick up the hexkey case and to extract the hexkey itself.
@@ D.4 Robot Hardware Spec
We conduct our real-world experiments using a custom-built robot named Anubis. The platform features a teleoperation system inspired by the Mobile ALOHA setup (Fu et al., 2024). Each arm has 6 DoF and is equipped with a parallel gripper and a wrist-mounted camera. At the center of the robot, an Intel RealSense camera is mounted on a height-adjustable mechanism, serving as the ego-centric view camera. Details are described in Table˜8. Anubis is equipped with a 3-wheel omni-directional base that supports planar locomotion; however, in this work, the mobility feature is not utilized.
@@ E.1 Tabletop-Sim
To test bimanual policies in simulation, we developed Tabletop-Sim, a new benchmark specifically engineered to evaluate dexterous manipulation skills, in contrast to other benchmarks (Mu et al., 2025) that primarily focus on task diversity. The code are publicly available at https://github.com/jellyho/Tabletop-Sim. The benchmark comprises four single-task environments and one multi-task setup. Our task selection was guided by the taxonomy in DexMimicGen (Jiang et al., 2025), which categorizes bimanual tasks into: (1) parallel (two arms are doing separate tasks simultaneously), (2) coordinated (two arms are closely working together), and (3) sequential (one arm completes the task, and the other arm takes over) interactions. Using a custom controller similar to GELLO (Wu et al., 2024), we collected 50 demonstrations for each single-task and 60 for the multi-task environment.
@@
Furthermore, to evaluate policy robustness, we established two difficulty settings for the four single-tasks. The original tasks are designated as the Easy setting, while a Hard variant for each task incorporates challenging variations such as different textures, object models, and the presence of distractor objects. Figure˜16 presents snapshots of each task.
@@ E.2 Quantitative Results
This section describes the detailed results for the simulation tasks. The results for Tabletop-Sim are listed in Table˜9, while the results for the RoboTwin 2.0 benchmark are in Table˜10. For RoboTwin, the results for other baselines were referenced from the official benchmark results.